OpenStudy (anonymous):

The equation of the lines joining the origin to the points of intersection of the line 3x + y = 12 intercepted between the axes of coordinates are

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

u mean lines or line

OpenStudy (anonymous):

Options: (1) y = 1/ 2 x and y = x ( 2) y = x and y = − x (3) y = 3/ 2 x and y = 6x (4) none of these

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think 4

OpenStudy (anonymous):

@experimentX

OpenStudy (ghazi):

the required equation is between the points (0,0) and (2,6) \[y-0=\frac{ 6-0 }{ 2-0 } (x-0)\] giving you y=3x you are done ;)

OpenStudy (anonymous):

lol wat is (2,6)

OpenStudy (ghazi):

2,6 is the intersection of line 3x+y=12 between the axes

OpenStudy (ghazi):

clear? lol

OpenStudy (anonymous):

nope..)

OpenStudy (ghazi):

@Yahoo! put x=0 and y=0 in the line 3x+y=12 to get those points

OpenStudy (ghazi):

i mean find the intercepts of the line dude

OpenStudy (anonymous):

3x+y=12 x-inter = 4 y-inter = 12 i got this...)

OpenStudy (anonymous):

but did nt get (2,6)

OpenStudy (ghazi):

yes and then do your x intercept is (4,0) and y is (0,6) are you kidding me.....man (2,6) is the point of intersection of the line 3x+y=12 THE MID POINT ...are you @Yahoo! lol

OpenStudy (ghazi):

i mean you get your intercept

OpenStudy (ghazi):

\[x=\frac{ 4+0 }{ 2 }=2 and y=\frac{ 12+0 }{ 2 }=6\] i chose mid point because your question says the point of intersection of line and that is the mid point

OpenStudy (ghazi):

@Yahoo! clear?

OpenStudy (anonymous):

Yup..Got it..) thxx..

OpenStudy (ghazi):

wow, you shocked me lol

OpenStudy (anonymous):